OpenStudy (anonymous):

(-5a + 7b -c) - (-8b +9c -8d)

OpenStudy (callisto):

(-5a + 7b -c) - (-8b +9c -8d) = (-5a + 7b -c) - (-8b) -(+9c) -(-8d) = -5a + 7b -c +8b -9c +8d = -5a +7b +8b -c -9c +8d = Can you do the last step?